Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 452     Entry time: 21 Nov 2023, 23:29
Author: Adam Round 
Group: SPB 
Subject: Shift end 

Staff: AR, MS, DM, RL, RdW

X-ray delivery

  • 16.9 keV, up to 30 pulses and 1.1 MHz for alignment, 1 pulse for data collection
  • stable delivery

Optical laser delivery

  • None

Achievements/Observations

  • Beam aligned down to zyla cam 3
  • CRLs were inserted but created too much speckles on Jungfrau and we couldn't reach the expected 80 um beam size for the users lenses (arms 6 and 7 unusable, arm 4 hole feature in beam)
  • user lenses aligned
  • User beamstop installed to block 1st order line in horizontal direction
  • Collected several speckle tracking data scans without our CRLs - to give initial estimates for focus
  • Jungfrau timing confirmed - signal observed

Issues

  • Could not see signal with the thin sample and the CRLs inserted
  • clear aberations from our CRLs
  • CRL arm 4 motors were not moving. To recover it, we connected the motor and encoder cables from the pinhole arm.
ELOG V3.1.4-7c3fd00